OpenStudy (anonymous):

Marge drove 348 miles in four and a half hours. At that rate, how many miles could she drive in 10 hours?

OpenStudy (anonymous):

348/4.5 = x/10 proportion so cross multiply

OpenStudy (anonymous):

773 thank you

OpenStudy (anonymous):

348 is equivalent to X; 4.5 is equivalent to 10; since the independent variable is the number of hours, that becomes the denominator: 348/4.5 = X/10 ; alternatively it can be written as X/10=348/4.5

OpenStudy (anonymous):

No problem.